OpenStudy (anonymous):

[\sqrt;[27[[x]^[10]]]]

OpenStudy (adamaero):

\[\sqrt{27x^{10}}\]

OpenStudy (adamaero):

3*9 = 27 (x^2)(x^2)(x^2)(x^2)(x^2) = x^10

OpenStudy (adamaero):

Figure it out from there...

OpenStudy (nurali):

\[\sqrt{3*3*3*x ^{10}}\]

OpenStudy (nurali):

\[3\sqrt{3x ^{10}}\]

OpenStudy (nurali):

\[3x^5\sqrt{3}\]

OpenStudy (nurali):

Welcome to Openstudy.